# Melbourne, VIC 3010, Australia

### v2

@@ 6.1 Spacetime filling branes
The general supersymmetry-preserving (or BPS) 11d background of coincident M2-branes can be written as
@@
and $\alpha_{2}=32\pi^{2}l_{P}^{6}$ in the standard normalization, $i=0,1,2$ the directions along the branes. The second part of the metric corresponds to the cone over a 7d Sasaki–Einstein space $L$, which we denote as $X_{\circ}=C(L)$, with metric as in (59). The transverse space $X_{\circ}$ to the M2-brane directions has a conical singularity at $r=0$, and is only smooth when $L$ is a sphere, S7, with the round metric. In the general case, $X_{\circ}$ is a non-compact Calabi–Yau four-fold singularity. Some toric examples of resolutions of such spaces were given in section 4. Importantly, the above background is a supersymmetric solution only in the presence of a non-vanishing three-form field,
@@
with $R_{L}=(\alpha_{2}\,q)^{1/6}$, and $\omega_{\mathrm{AdS}_{4}}$ the volume form with unit radius. We can canonically define the integer $N$ as the M2-brane electric charge (here at leading order as we consider the two derivative theory, $N_{\rm M2}=N$),
@@
To clarify the holographic picture, schematically shown in Figure 2, the AdS/CFT correspondence applies precisely in the near-horizon limit, $\mathrm{AdS}_{4}\times L$, and relates M-theory on this background to the gauge theory localized on the M2-branes.  [^21] While we focus on M2-branes, the discussion can almost directly apply to D3 or M5-branes, with only minor differences due to their dimensionalities. In the absence of a non-perturbative definition of the bulk theory—which we argue is captured by equivariant topological strings—the low-energy supergravity description is typically applied in the bulk and related to a strong coupling limit in the dual field theory. 11d supergravity can be further reduced on the compact space $L$ to an effective four-dimensional gauged supergravity, in a typical example of the so called flux compactification in string/M-theory. In this case, the effective Newton constant in the lower dimensional theory, $G_{\mathrm{4d}}$, depends explicitly on the internal volume, see e.g. (Marino:2011nm, , Section 5),
@@
with the combination on the left hand side setting the overall scale of 4d supergravity quantities (since $R_{L}$ is proportional to the $\mathrm{AdS}_{4}$ radius). Notice that the M2-brane charge $N$ is a natural parameter both in the field theory description, corresponding to the gauge group rank, and in the supergravity description, via (205). On the other hand, the geometry of $L$ itself only depends on the flux $N$ in its overall size and not topologically. This is the fundamental origin of the need for different ensembles between the topological string ($\lambda$) and the field theory/supergravity ($N$ or $G_{\mathrm{N}}$) descriptions.
@@
with an implicit summation over repeated indices. We thus see that the Reeb vector on $X$ actually parametrizes the standard torus action, such that $\epsilon_{i}$ are the equivariant parameters in the sense of Section 2. In the case when we keep the 4d space maximally symmetric (i.e. in Euclidean signature pure $\mathrm{AdS}_{4}$ is sliced into round S3 coordinates),  [^22] In fact, we can also relax the assumption on the boundary metric without changing the internal topology by considering a squashed S3 boundary that only preserves $U(1)\times U(1)$ isometry, see Martelli:2011fu . This introduces an additional deformation parameter, usually denoted by $b$, which also generalizes the supersymmetric constraint. We elaborate further on these issues in Cassia:2024 . supersymmetry imposes a particular condition on these equivariant parameters,
@@
with a Kähler form $\omega_{B}$ on the Kähler metric transverse to the Reeb foliation, see Couzens:2018wnk ; Gauntlett:2018dpc ; Hosseini:2019use ; Hosseini:2019ddy ; Gauntlett:2019roi for details and Katmadas:2015ima for a different approach to this construction. We can thus view the space $L$ as a circle fibration over a compact space $B$ of three complex dimensions. This gives a more direct geometric meaning of the $\lambda$ parameters as responsible for the physical (non-equivariant) size of the compact space $B$, cf. the discussion around (74).
Given this setup, there are many geometric relations that have been shown to hold between the Sasakian manifold $L$, the Kähler space $B$, and the cone $X_{\circ}$, see again Section 2.3 and Martelli:2005tp ; Butti:2005vn ; Butti:2005ps ; Martelli:2006yb ; Amariti:2011uw ; Couzens:2018wnk ; Gauntlett:2018dpc ; Hosseini:2019use ; Hosseini:2019ddy ; Gauntlett:2019roi ; Gauntlett:2019pqg ; Boido:2022mbe . A quantity of particular interest is the Sasakian volume $\mathrm{vol}_{L}(\epsilon)$, which now depends on the parameters $\epsilon_{i}$ [^23] Note that the Sasakian volume $\mathrm{vol}_{L}(\epsilon)$ should be thought of as an off-shell quantity and is only related to the quantity entering Eq’s (204) and (205) after a suitable extremization fixing the parameters $\epsilon$ on their on-shell values. Its relation to the equivariant volume was first shown in Martelli:2006yb (passing through the equivariant index) and revisited in Martelli:2023oqk ,
$\mathrm{vol}_{L}(\epsilon)=\frac{2\pi^{4}}{3!}\,\mathrm{vol}_{X}(t=0,\epsilon)\ .$ (209)
This relation holds at vanishing Kähler parameters corresponding to the mesonic twist describing the conical metric of $X_{\circ}$, (59).
@@
The master volume plays a crucial role in holography, as it directly relates to the supergravity action and in turn to the dual holographic free energy. From the present point of view, the most important insight on the relation between $\lambda$ and $N$ comes from a holographic argument proposed in Couzens:2018wnk ; Gauntlett:2018dpc ; Hosseini:2019use ; Hosseini:2019ddy ; Gauntlett:2019roi . The R-charges of the dual field theory operators associated to wrapping M5-branes on toric divisors $T_{i}\in L$ are given by
@@
where the $\epsilon_{i}$ are subject to the mesonic constraint, (71). On the field theory side, in the large $N$ limit corresponding to the two derivative supergravity approximation we took here, the R-charges are linear in $N$,
@@
where supersymmetry further requires the condition
@@
One can now identify the rescaled R-charges $\Delta_{i}$ of the M2-brane theory with the equivariant parameters $\epsilon_{i}$, which we carry out explicitly in Cassia:2024  [^24] While we focused on an off-shell point of view here, the same setup allows also a description within on-shell 4d supergravity. In this picture the generalization to arbitrary values of $\epsilon_{i}$ corresponds to switching on scalar profiles in 4d and leads to a more general solution that no longer preserves all $\mathrm{AdS}_{4}$ symmetries, see Freedman:2013oja . This is also reflected in the fact that a generic value of the $\Delta_{i}$ deformations, corresponding to R-charge assignments of the dual field theory on round S3, breaks the conformal symmetry. The important point here is the remaining equation that relates the first derivative of the genus-zero constant maps to the M2-brane charge $N$,
@@
This is precisely the Legendre transform depicted schematically in Figures 1 and 2 relating the topological string picture to the effective supergravity and dual field theory. The extremization holds in the limit of two derivative supergravity and corresponding large $N$ field theory limit, corresponding to the first approximation of a full integral transform between the $\lambda$ and $N$ ensembles. We devote our companion paper, Cassia:2024 , to the extension of this relation to finite $N$ in field theory and higher derivatives in supergravity, based on the additional genus-one constant maps term we have defined here. This analysis is greatly facilitated by the host of important results on the exact M2-brane partition function on round and squashed S3, see Kapustin:2009kz ; Hama:2011ea ; Imamura:2011wg ; Fuji:2011km ; Marino:2011eh ; Nosaka:2015iiw ; Hatsuda:2016uqa ; Chester:2021gdw and references therein.
@@ 6.2 Wrapped branes
i.e. the total space of the fibration of $L$ over $\Sigma$. We leave unspecified the warp factor $\sigma$ in the metric, also a function of the internal space, as it does not affect the topology. Similarly to (204) we impose the flux quantization condition (still only at two derivative level),
@@
where $S_{i}\hookrightarrow M$ are submanifolds representatives of the generators of the free part of $H_{7}(M,\mathbb{Z})$. Moreover, the generators $S_{i}$ are related to the “divisors” $T_{i}\hookrightarrow L$, introduced in the previous section, as they fit into the diagram
@@
From a lower-dimensional perspective, if we require $\Sigma$ to be toric, these backgrounds correspond to supersymmetric rotating black holes with spherical or spindle horizon topology, $\Sigma=\mathbb{WP}^{1}_{a,b}$, as discussed in Cacciatori:2009iz ; Halmagyi:2013sla ; Hristov:2018spe ; Hristov:2019mqp ; Bobev:2020pjk ; Ferrero:2020twa ; Couzens:2021cpk ; Ferrero:2021etw ; Hristov:2023rel . Similar to the discussion surrounding (205), the lower-dimensional Newton constant arises from explicit integration over the internal space. In the context of black hole physics, the two-dimensional Newton constant also plays a key role, being related to the Bekenstein–Hawking entropy of the black holes, $S_{\text{BH}}=1/(4G_{\mathrm{2d}})$. Again, this M2-brane construction admits a dual holographic description via localization techniques, as explored in Benini:2015noa ; Benini:2015eyy ; Hosseini:2022vho ; Bobev:2023lkx ; Bobev:2024mqw ; Inglese:2023wky ; Colombo:2024mts and references therein, which we revisit in Cassia:2024 .
@@
with $\epsilon_{i}$ and $n_{i}$ obeying the mesonic constraints on the fiber $X$, (71) and (194), respectively. The above equations correspond holographically to $\mathcal{I}$-extremization, Benini:2015eyy . It would be interesting to understand the holographic role of the mesonic twist on $Y$ (on top of the one on $X$ imposed above), which has not been explored in Martelli:2023oqk . We come back to this point in Cassia:2024 .
In this section, we have provided a basic overview of flux compactifications and their conceptual connection to the equivariant topological string on toric Calabi–Yau cones. While additional subtleties and details are involved in establishing the precise relation to the M2-brane partition functions, as outlined, these are explored in greater depth in our companion paper, Cassia:2024 . For the remainder of this work, we focus on the special and intriguing case of vanishing flux, which bypasses many of these complexities and, in a sense, completes the examples introduced in Section 5.
7 Vanishing flux
In the previous section, we discussed background solutions with coincident branes in 11d supergravity and their near-horizon limit, leading to flux compactifications. Another case with vanishing fluxes involves branes that either entirely wrap cycles of the internal space or are localized at isolated points.  [^26] We are very heuristic in the description of the brane construction here since there exist many alternative descriptions in different duality frames, see e.g. Maldacena:1997de for further discussion. For our present purposes we only need to know the topology of the internal space from 11d point of view, the specific brane construction is immaterial. From a low-energy perspective, these configurations result in asymptotically flat 4d spacetime, either empty or populated by black hole centers. In this case, the anomaly cancellation condition ensures no overall flux through the compact manifold Sethi:1996es ; Witten:1996md . The equations of motion require the internal space to be Ricci flat, with unfixed Kähler moduli. As a result, the effective Newton constant is independent of the internal space volume.

### v4

@@ 6.1 Spacetime filling branes
One type of supersymmetry-preserving (or BPS) 11d background of multiple spacetime filling M2-branes can be written as  [^24] There are even more general backgrounds, which we have neglected in the presented form of the metric and three-form backgrounds, which correspond to further deformations of the underlying M2-brane theory, see e.g. Pope:2003jp ; Bena:2004jw and references thereof. Even though we do not consider the explicit 11d solutions, the topological string description still allows to identify such deformations in the dual field theory picture.
@@
and $\alpha_{2}=32\pi^{2}l_{P}^{6}$ in the standard normalization. Here the index $i=0,1,2$ labels the coordinates along the spacetime directions where the branes are located. The second part of the metric corresponds to the cone over a 7d Sasaki–Einstein space $L$, which we denote as $X_{\circ}=C(L)$, with metric as in (59). The transverse space $X_{\circ}$ to the M2-brane directions has a conical singularity at $r=0$, and is only smooth when $L$ is a sphere, S7, with the round metric. In the general case, $X_{\circ}$ is a non-compact Calabi–Yau four-fold singularity. Some toric examples of resolutions of such spaces were given in section 4. Importantly, the above background is a supersymmetric solution only in the presence of a non-vanishing three-form field,
@@
with $R_{L}=(\alpha_{2}\,q)^{1/6}$, and $\omega_{\mathrm{AdS}_{4}}$ the volume form with unit radius. We can canonically define the integer $N$ as the M2-brane electric charge (here at leading order as we consider the two derivative theory, $N_{\rm M2}=N$, see App. A),
@@
To clarify the holographic picture, schematically shown in Figure 2, the AdS/CFT correspondence applies precisely in the near-horizon limit, $\mathrm{AdS}_{4}\times L$, and relates M-theory on this background to the gauge theory localized on the M2-branes.  [^25] While we focus on M2-branes, the discussion can almost directly apply to D3 or M5-branes, with only minor differences due to their dimensionalities. In the absence of a non-perturbative definition of the bulk theory—which we argue is captured by equivariant topological strings—the low-energy supergravity description is typically applied in the bulk and related to a strong coupling limit in the dual field theory. 11d supergravity typically allows for consistent truncations of finite sets of KK modes on the compact space $L$ to an effective four-dimensional gauged supergravity, see deWit:1986oxb ; Gauntlett:2007ma ; Gauntlett:2009zw ; Cassani:2011fu ; Nicolai:2011cy ; Cassani:2012pj . In this case, the effective Newton constant in the lower dimensional theory, $G_{\mathrm{4d}}$, depends explicitly on the internal volume, see e.g. (Marino:2011nm, , Section 5),
@@
with the combination on the left hand side setting the overall scale of 4d supergravity quantities (since $R_{L}$ is proportional to the $\mathrm{AdS}_{4}$ radius).  [^26] In presence of higher-derivative corrections in 11d, we expext that the above equation holds with the substitution of $N$ with $N_{\rm M2}$, see App. A. Notice that the M2-brane charge $N$ is a natural parameter both in the field theory description, corresponding to the gauge group rank, and in the supergravity description, via (205). On the other hand, the geometry of $L$ itself only depends on the flux $N$ in its overall size and not topologically. This is the fundamental origin of the need for different ensembles between the topological string ($\lambda$) and the field theory/supergravity ($N$ or $G_{\mathrm{N}}$) descriptions.
@@
with an implicit summation over repeated indices. We thus see that the Reeb vector on $X$ actually parametrizes the standard torus action, such that $\epsilon_{i}$ are the equivariant parameters in the sense of Section 2. In the case when we keep the 4d space maximally symmetric (i.e. in Euclidean signature pure $\mathrm{AdS}_{4}$ is sliced into round S3 coordinates),  [^27] In fact, we can also relax the assumption on the boundary metric without changing the internal topology by considering a squashed S3 boundary that only preserves $U(1)\times U(1)$ isometry, see Martelli:2011fu . This introduces an additional deformation parameter, usually denoted by $b$, which also generalizes the supersymmetric constraint. We elaborate further on these issues in Cassia:2025jkr . supersymmetry imposes a particular condition on these equivariant parameters,
@@
with a Kähler form $\omega_{B}=\mathrm{d}\eta$ transverse to the Reeb foliation, see Couzens:2018wnk ; Gauntlett:2018dpc ; Hosseini:2019use ; Hosseini:2019ddy ; Gauntlett:2019roi for details and Katmadas:2015ima for a different approach to this construction. We can thus view the space $L$ as a circle fibration over a compact space $B$ of three complex dimensions. This gives a more direct geometric meaning of the $\lambda$ parameters as responsible for the physical (non-equivariant) size of the compact space $B$, cf. the discussion around (74).
In what follows, we will also need to consider certain submanifolds of $L$. Assuming that the cone over $L$ is indeed a toric CY, we can deduce that to each toric divisor $D_{i}$ of $C(L)$, i.e. the locus where the $i$-th homogeneous coordinate vanishes, one can associate a corresponding (real) codimension-2 submanifold $T_{i}$ in $L$, obtained as the intersection of the toric divisor and the base of the cone,
$T_{i}:=D_{i}\cap L\,.$ (209)
By analogy with the toric case, we will refer to the $T_{i}$’s as Sasakian divisors of $L$.
Given this setup, there are many geometric relations that have been shown to hold between the Sasakian manifold $L$, the Kähler space $B$, and the cone $X_{\circ}$, see again Section 2.3 and Martelli:2005tp ; Butti:2005vn ; Butti:2005ps ; Martelli:2006yb ; Amariti:2011uw ; Couzens:2018wnk ; Gauntlett:2018dpc ; Hosseini:2019use ; Hosseini:2019ddy ; Gauntlett:2019roi ; Gauntlett:2019pqg ; Boido:2022mbe . A quantity of particular interest is the Sasakian volume $\mathrm{vol}_{L}(\epsilon)$, which now depends on the parameters $\epsilon_{i}$.  [^28] Note that the Sasakian volume $\mathrm{vol}_{L}(\epsilon)$ should be thought of as an off-shell quantity and is only related to the quantity entering Eq’s (204) and (205) after a suitable extremization fixing the parameters $\epsilon$ on their on-shell values. Its relation to the equivariant volume was first shown in Martelli:2006yb (passing through the equivariant index) and revisited in Martelli:2023oqk ,  [^29] Similarly to the definition of the equivariant volume, the Sasakian volume here is normalized differently with respect to Martelli:2006yb ; Martelli:2023oqk and the related references.
$\mathrm{vol}_{L}(\epsilon)=\mathrm{vol}_{X}(t=0,\epsilon)\ ,$ (210)
where we remark that, in our conventions, the equivariant parameters $\epsilon_{i}$ are redundant, while the original definition of the Sasakian volume only depends on the effective equivariant parameters (they enter the computation of the Sasakian volume as coefficients of the Reeb vector on the basis of abelian isometries of $L$). In order to match the parametrizations of the volumes, we make use of the fact that the Sasakian volume can be obtained as the cohomological limit of the K-theoretic equivariant index, which we can naturally upgrade to a function of all equivariant parameters, including the non-effective ones. In this way, we can regard (210) as a definition of this upgraded function, which we obtain as the cohomological limit of the fully equivariant index. In this way, the volume $\mathrm{vol}_{L}(\epsilon)$ reduces by construction to the usual definition of Sasakian volume once we lift the redundancy in the $\epsilon$-parametrization.
We notice, moreover, that a similar definition of volume can be formulated for the Sasakian divisors $T_{i}$. In fact, since they arise as intersections between toric divisors in $X$ and the base $L$, one can define their volume as
$\mathrm{vol}_{T_{i}}(\epsilon)=\mathrm{vol}_{D_{i}}(t=0,\epsilon)\ ,$ (211)
which is equivalent to (210), once we substitute $T_{i}$ and $D_{i}$ for $L$ and $X$, respectively. These relations, (210)-(211), already establish the link between the internal volume extremization program initiated in Martelli:2006yb and the present approach.
@@
The master volume plays a crucial role in holography, as it directly relates to the supergravity action and in turn to the dual holographic free energy. From the present point of view, the most important insight on the relation between $\lambda$ and $N$ comes from a holographic argument proposed in Couzens:2018wnk ; Gauntlett:2018dpc ; Hosseini:2019use ; Hosseini:2019ddy ; Gauntlett:2019roi which concerns certain supersymmetric submanifolds of $L$, which are identified with the Sasakian divisors (209). Holography then predicts that the R-charges of the dual field theory operators associated to wrapping M5-branes on the $T_{i}$ are given by
@@
where the $\epsilon_{i}$ are subject to the mesonic constraint, (71). On the field theory side, in the large $N$ limit corresponding to the two derivative supergravity approximation which we took here, the R-charges are linear in $N$,
@@
where supersymmetry further imposes the condition
@@
One can now identify the rescaled R-charges $\Delta_{i}$ of the M2-brane theory with the equivariant parameters $\epsilon_{i}$, which we carry out explicitly in Cassia:2025jkr .  [^30] While we focused on an off-shell point of view here, the same setup allows also a description within on-shell 4d supergravity. In this picture the generalization to arbitrary values of $\epsilon_{i}$ corresponds to switching on scalar profiles in 4d and leads to a more general solution that no longer preserves all $\mathrm{AdS}_{4}$ symmetries, see Freedman:2013oja . This is also reflected in the fact that a generic value of the $\Delta_{i}$ deformations, corresponding to R-charge assignments of the dual field theory on round S3, breaks the conformal symmetry. The important point here is the remaining equation that relates the first derivative of the genus-zero constant maps to the M2-brane charge $N$,
@@
This is precisely the Legendre transform depicted schematically in Figures 1 and 2 relating the topological string picture to the effective supergravity and dual field theory. The extremization holds in the limit of two derivative supergravity and corresponding large $N$ field theory limit, corresponding to the first approximation of a full integral transform between the $\lambda$ and $N$ ensembles. We devote our companion paper, Cassia:2025jkr , to the extension of this relation to finite $N$ in field theory and higher derivatives in supergravity, based on the additional genus-one constant maps term we have defined here. This analysis is greatly facilitated by the host of important results on the exact M2-brane partition function on round and squashed S3, see Kapustin:2009kz ; Hama:2011ea ; Imamura:2011wg ; Fuji:2011km ; Marino:2011eh ; Nosaka:2015iiw ; Hatsuda:2016uqa ; Chester:2021gdw and references therein.
@@ 6.2 Wrapped branes
i.e. the total space of the fibration of $L$ over $\Sigma$. We leave unspecified the warp factor $\sigma$ in the metric, also a function of the internal space, as it does not affect the topology. Similarly to (204) we impose the flux quantization condition (still only at two derivative level, $N_{{\rm M2},i}=n_{i}\,N$, see App. A),
@@
where $S_{i}\hookrightarrow M$ are submanifolds representatives of the generators of the free part of $H_{7}(M,\mathbb{Z})$. Moreover, the generators $S_{i}$ are related to the Sasakian divisors $T_{i}\hookrightarrow L$, introduced in (209), as they fit into the diagram
@@
From a lower-dimensional perspective, if we require $\Sigma$ to be toric, these backgrounds correspond to supersymmetric rotating black holes with spherical or spindle horizon topology, $\Sigma=\mathbb{WP}^{1}_{a,b}$, as discussed in Cacciatori:2009iz ; Halmagyi:2013sla ; Hristov:2018spe ; Hristov:2019mqp ; Bobev:2020pjk ; Ferrero:2020twa ; Couzens:2021cpk ; Ferrero:2021etw ; Hristov:2023rel . Similar to the discussion surrounding (205), the lower-dimensional Newton constant arises from explicit integration over the internal space. In the context of black hole physics, the two-dimensional Newton constant also plays a key role, being related to the Bekenstein–Hawking entropy of the black holes, $S_{\text{BH}}=1/(4G_{\mathrm{2d}})$. Again, this M2-brane construction admits a dual holographic description via localization techniques, as explored in Benini:2015noa ; Benini:2015eyy ; Hosseini:2022vho ; Bobev:2023lkx ; Bobev:2024mqw ; Inglese:2023wky ; Colombo:2024mts and references therein, which we revisit in Cassia:2025jkr .
@@
with $\epsilon_{i}$ and $n_{i}$ obeying the mesonic constraints on the fiber $X$, (71) and (194), respectively. The above equations correspond holographically to $\mathcal{I}$-extremization, Benini:2015eyy . It would be interesting to understand the holographic role of the mesonic twist on $Y$ (on top of the one on $X$ imposed above), which has not been explored in Martelli:2023oqk . We come back to this point in Cassia:2025jkr .
In this section, we have provided a basic overview of M2-brane backgrounds and their conceptual connection to the equivariant topological string on toric Calabi–Yau cones. While additional subtleties and details are involved in establishing the precise relation to the M2-brane partition functions, as outlined, these are explored in greater depth in our companion paper, Cassia:2025jkr . For the remainder of this work, we focus on the special and intriguing case of vanishing fluxes, which bypasses many of these complexities and, in a sense, completes the examples introduced in Section 5.
7 Vanishing flux backgrounds
In the previous section, we discussed background solutions with coincident branes in 11d supergravity, exhibiting non-trivial flux. The special case of vanishing fluxes that we examine now involves branes that either entirely wrap cycles of the internal space or are localized at isolated points.  [^32] We are very heuristic in the description of the brane construction here since there exist many alternative descriptions in different duality frames, see e.g. Maldacena:1997de for further discussion. For our present purposes we only need to know the topology of the internal space from 11d point of view, the specific brane construction is immaterial. From a low-energy perspective, these configurations result in asymptotically flat 4d spacetime, either empty or populated by black hole centers. In this case, the anomaly cancellation condition ensures no overall flux through the compact manifold Sethi:1996es ; Witten:1996md . The equations of motion require the internal space to be Ricci flat, with unfixed Kähler moduli. As a result, the effective Newton constant is independent of the internal space volume.